An electric power station that operates at 15 kV and uses a 10:1 step-up ideal transformer is producing 360 MW (Mega-Watt) of power that is to be sent to a big city which is located 260 km away with only 2.0% loss. Each of the two wires are made of copper (resistivity = 1.68 × 10^ -8 Ω.m). What is the diameter of the wires?
Answer #1

From Transformer turns ratio

VS/VP=NS/NP

VS/15 =(10/1)

Vs=150 KV (transmitted voltage)

Current transmitted

I=P/Vs=(360*106)/(150*103)=2400 A

Given Transmission line losses

Ploss=0.02*360 =7.2 MW

Since Ploss=I2R

24002*R =7.2*106

R=1.25 ohms

Since Resistance is given by

R=pL/A

=>A=pL/R =(1.68*10-8)(2*260*103)/1.25

A=6.9888*10-3 m2

A=pi*D2/4=6.9888*10-3

D=0.09433 m =9.433 cm